# RePO: ReLU-based Preference Optimization

### v1

@@ Abstract
Aligning large language models (LLMs) with human preferences is critical for real-world deployment, yet existing methods like RLHF face computational and stability challenges. While DPO establishes an offline paradigm with single hyperparameter $\beta$, subsequent methods like SimPO reintroduce complexity through dual parameters ($\beta$, $\gamma$). We propose ReLU-based Preference Optimization (RePO), a streamlined algorithm that eliminates $\beta$ via two advances: (1) retaining SimPO’s reference-free margins but removing $\beta$ through gradient analysis, and (2) adopting a ReLU-based max-margin loss that naturally filters trivial pairs. Theoretically, RePO is characterized as SimPO’s limiting case ($\beta\to\infty$), where the logistic weighting collapses to binary thresholding, forming a convex envelope of the 0-1 loss. Empirical results on AlpacaEval 2 and Arena-Hard show that RePO outperforms DPO and SimPO across multiple base models, requiring only one hyperparameter to tune.
LLM Alignment, RLHF, Data Selection
@@ 1 Introduction
Aligning large language models (LLMs) with human preferences is essential for their effective deployment in real-world applications (Anil et al., 2023; Touvron et al., 2023; OpenAI, 2023; Bubeck et al., 2023), ensuring that their outputs are beneficial, accurate, and adhere to human values (Ouyang et al., 2022) while minimizing potential safety risks. A primary approach for achieving this alignment is Reinforcement Learning from Human Feedback (RLHF). It typically entails a multi-stage process: initially, a reward model is trained using ranked preference data; subsequently, the LLM is optimized with the guidance of reward model, using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017). Despite its effectiveness, RLHF faces significant challenges, such as high computational costs and training instability (Rafailov et al., 2023; Zhao et al., 2023).
Recent research has been exploring simpler offline alternatives. Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one such approach. It reparameterizes the reward function in RLHF, bypassing the explicit learning of the reward model, and thus enabling direct training of LLMs with human preference data. Formally, the DPO objective is a logistic-log loss, which applies a log-sigmoid activation to the reward margin between the preferred and less-preferred responses to the same input. Here, the reward is derived from comparing the policy model of the LLM with a reference model. Building on this, follow-up studies, especially SLiC-HF (Zhao et al., 2023) and SimPO (Meng et al., 2024), focus on further simplifications on the loss function and reward margins separately. Specifically, SLiC-HF replaces the logistic-log loss in DPO with a hinge loss, achieving a faster-decaying tail. SimPO, on the other hand, introduces reference-free reward margins with logistic-log loss, offering a more streamlined and efficient approach. While advancing the field and achieving leading performance, they introduce additional complexity on hyperparameter tuning. SimPO inherits DPO’s reward margin strength controller $\beta$ and adds a target reward margin $\gamma$. Similarly, SLiC-HF utilizes $\gamma$, but also introduces a regularization weight $\lambda$. The necessity of dual-tuning these hyperparameters poses significant challenges in model training and deployment. This leads to a natural question:
Can we develop a simpler offline preference optimization algorithm?
Figure 1: RePO, SimPO, and DPO primarily differ in their loss functions, as highlighted in the shaded box. RePO either outperforms or achieves comparable performance to DPO and SimPO across various settings on AlpacaEval 2, with only a single hyperparameter $\gamma$. “IT” denotes pre-trained instruction-tuned models. e.g., “Gemma2-IT 9B” refers to Gemma2-9B (instruct setup).
In this paper, we propose a simple idea: simultaneously simplifying the log-sigmoid activation and reward margins, to achieve comparable or even better performance. For reward margins, we adopt the reference-free approach of SimPO but eliminate the hyperparameter $\beta$. For the log-sigmoid activation, we employ ReLU to the reference-free reward margins, following a max-margin (support vector machine) paradigm (Boser et al., 1992; Cortes & Vapnik, 1995), similar to SLiC-HF. Consequently, our method can be interpreted as “SimPO without $\beta$” and “SLiC-HF without $\lambda$”. Interestingly, it bridges these methods by omitting a core component from each while maintaining competitive performance, demonstrating the effectiveness of this streamlined approach. We term this ReLU-based Preference Optimization RePO.
Theoretically, we demonstrate that RePO emerges as the natural endpoint of SimPO’s architectural evolution: when $\beta\to\infty$, SimPO’s logistic weighting collapses to RePO’s binary thresholding (Lemma 3.1). This asymptotic behavior motivates our key theoretical insight — the ReLU loss in RePO is precisely the convex envelope of the 0-1 loss (Theorem 4.2), inheriting its global optimality while enabling tractable gradient-based optimization. Practically, this formulation provides the following advantages:
  • Simplified Hyperparameter Tuning: RePO removes the hyperparameter $\beta$, requiring only the tuning of $\gamma$ within the range $[0,1]$. Empirical results show that RePO can achieve performance comparable to SimPO on three models (Llama, Mistral, and Gemma) even with a fixed $\gamma=0.5$, significantly reducing the complexity of hyperparameter tuning.
  • Effective Data Filtering: By using ReLU instead of a continuous sigmoid function, RePO interprets the target reward margin $\gamma$ as a strict “cutoff point”. This mechanism zeroes out gradients for preference pairs whose reward margin exceeds $\gamma$ (well-separated pairs), while assigning unit weight to less-separated pairs. This selective approach minimizes overfitting by focusing optimization on challenging data points.
  • Controllable Over-Optimization: The “cutoff point” $\gamma$ controls the mean of reward margin within a batch, representing a new metric for evaluating over-optimization. Experimental results suggest that this metric correlates with model behavior and can replace KL divergence as a simpler alternative, potentially controlling optimization without requiring a reference model.
While RePO effectively filters out well-separated pairs by zeroing their gradients, it assigns a unit gradient weight to all less-separated pairs. This implicit uniform weighting overlooks the varying importance of these pairs, potentially limiting its ability to prioritize more informative examples. To address this limitation, we extend the concept of RePO within SimPO and propose a new method, RePO++, by leveraging SimPO’s logistic-log loss to differentiate the importance of less-separated pairs. This approach assigns greater weights to less-separated pairs, while continuing to filter out well-separated pairs. Notably, only a single line of code is required to seamlessly integrate RePO into SimPO, resulting in RePO++. In a nutshell, our main contribution is ReLU-based Preference Optimization (RePO), a simple but effective reference-free algorithm for training language models from preferences. With its straightforward motivation, RePO is grounded in a principled theoretical framework and consistently outperforms DPO and SimPO across three base model settings (Mistral2-7B, Llama3-8B, and Gemma2-9B) on AlpacaEval 2 (cf. Figure 1). We hope our exploration will motivate people to rethink the fundamental roles of loss function design for preference optimization.
@@ 2 Preliminaries
Offline Alignment. In the offline alignment setting, we utilize a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where each entry consists of a prompt $x$ accompanied by two responses: $y_{w}$ is the preferred (winning) response, and $y_{l}$ is the less preferred (losing) response, both generated by a reference policy $\pi_{\text{ref}}$. The primary challenge is to find a latent reward function $r^{*}(x,y)$ that governs these preference relationships. Although $r^{*}(x,y)$ is not directly observable, the Bradley-Terry (BT) model (Bradley & Terry, 1952) provides a probabilistic framework for modeling the probability of $y_{w}$ over $y_{l}$ given $x$:
$\mathbb{P}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r^{*}(x,y_{w}))}{\exp(r^{*}(x,y_{w}))+\exp(r^{*}(x,y_{l}))}.$ (1)
Here, $r^{*}(x,y)$ represents the latent reward for a response $y$ in the context of prompt $x$. The aim of offline alignment is to develop a policy $\pi_{\theta}$ capable of accurately estimating this latent reward function using the dataset $\mathcal{D}$.
Reinforcement Learning from Human Feedback (RLHF). Traditional offline alignment methods often leverage reinforcement learning augmented with a KL-divergence regularization term in the reward objective (Bai et al., 2022; Ziegler et al., 2019; Ouyang et al., 2022). The optimization objective, parameterized by a regularization parameter $\beta>0$, is expressed as:
$\mathop{\max}_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y\mid x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)),$ (2)
where $r_{\phi}(x,y)$ represents the reward function trained using the BT model on the preference data, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the fixed reference policy, typically derived from supervised fine-tuning. The KL-divergence term ensures that $\pi_{\theta}$ remains close to $\pi_{\text{ref}}$, mitigating risks of excessive divergence during training.
Directed Preference Optimization (DPO). DPO (Rafailov et al., 2023) stands out as a leading method for offline preference optimization by eliminating the need for an explicit reward model. Instead, it reformulates the reward function $r(x,y)$ as a closed-form expression based on policy ratios:
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (4)
$\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right)\right],$
@@
Sequence Likelihood Calibration (SLiC-HF). SLiC-HF (Zhao et al., 2023) advances preference optimization with two key innovations: (1) it employs a sequence-level calibration loss that contrasts the log-probability difference between preferred and dispreferred responses using a margin $\gamma$, and (2) it integrates a regularization term to prevent divergence from the SFT policy, avoiding the need for an explicit KL penalty. The SLiC-HF loss function is defined as:
$\mathcal{L}_{\text{SLiC-HF}}(\pi_{\theta})=$ $\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\Big{[}\text{ReLU}\Big{(}-\Big{(}\log{\pi_{\theta}(y_{w}\mid x)}$ (5)
$-\log\pi_{\theta}(y_{l}\mid x)-\gamma\Big{)}\Big{)}-\lambda\log{\pi_{\theta}(y_{w}\mid x)}\Big{]}.$
Simple Preference Optimization (SimPO). SimPO (Meng et al., 2024) advances preference optimization with two key innovations: (1) it normalizes the reward by the length of the response, calculating the average log-probability per token for a response under the policy $\pi_{\theta}$, and (2) it incorporates a target reward margin $\gamma$ to ensure that the reward difference between the preferred and less preferred responses exceeds this margin. The SimPO loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (6)
$\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|}-\gamma\right)\right)\right],$
@@
3 Analysis on RePO
In this section, we first present the expression for RePO and compare its performance with that of SimPO through experiments. We then analyze the advantages and key factors behind the success of RePO from the perspectives of gradient behavior, data filtering, and over-optimization.
3.1 RePO
Method Formulation. We propose RePO by simplifying the SimPO loss function through two key modifications: (i) removing the hyperparameter $\beta$, and (ii) replacing the log-sigmoid function with a ReLU activation. First, we define the implicit reward margin $M_{\theta}$ as:
@@ 3.1 RePO
which measures the policy’s relative preference between responses. Using $M_{\theta}$, the proposed loss function is:
@@
Comparison between RePO and SimPO / SLiC-HF. As formalized in Equation 8, RePO simplifies the SimPO framework by replacing the log-sigmoid activation with a ReLU operator and removing the need to tune the hyperparameter $\beta$. Simultaneously, it eliminates the explicit regularization term in SLiC-HF, thereby removing the associated hyperparameter $\lambda$. This dual simplification makes RePO a unified specialization of both SimPO and SLiC-HF, reducing hyperparameter tuning complexity while maintaining competitive performance, as demonstrated in Section 3.2.
Training Dynamics Monitoring. We track optimization progress with two metrics:
  • Batch-level reward margin:
    $m_{\text{batch}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{B}}[M_{\theta}],$ (9)
    which reflects the average reward margin within each training batch $\mathcal{B}$.
  • Dataset-level reward margin:
    $m_{\mathcal{D}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[M_{\theta}],$ (10)
    which measures the overall reward margin on the entire training set $\mathcal{D}$.
Here, $m_{\text{batch}}$ provides immediate feedback during each update, while $m_{\mathcal{D}}$ captures the global preference strength of the policy over the complete dataset.
Gradient Analysis. We analyze the gradient dynamics of RePO and SimPO to highlight their fundamental differences, the gradients are as follows:
@@
where $s_{\theta}=\sigma(\beta(-M_{\theta}+\gamma))$ is SimPO’s sigmoid weighting function. The terms $\nabla_{\theta,y_{w}}=\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)$ and $\nabla_{\theta,y_{l}}=\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)$ correspond to the gradients that increase the probability of the “winning” response $y_{w}$ and decrease the probability of the “losing” response $y_{l}$, respectively. The scaling factor $\beta$ in Equation 11 linearly amplifies gradient magnitudes but does not alter the relative update directions in adaptive optimizers like Adam (Kingma & Ba, 2014), as the momentum terms automatically normalize scale variations. We therefore omit $\beta$ in Figure 2 for clearer visualization of the weighting function shapes.
RePO’s gradient (Equation 12) applies uniform updates to pairs with $M_{\theta}<\gamma$, while SimPO (Equation 11) employs $\beta$-scaled continuous weighting through $s_{\theta}$. Figure 2 visualizes this contrast, showing how RePO’s binary thresholding emerges as a limiting case of SimPO when $\beta\to\infty$.
Figure 2: Gradient weighting functions of SimPO ($s_{\theta}$) and RePO ($\mathbb{I}(M_{\theta}<\gamma)$). As $\beta\to\infty$, $s_{\theta}$ converges to the binary indicator (dashed line), establishing RePO as the limit case of SimPO.
@@
Please check Appendix for all proofs. Lemma 3.1 establishes RePO as the asymptotic limit of SimPO with large $\beta$. This explains why RePO achieves comparable performance to SimPO (Section 3.2) while eliminating $\beta$ tuning. The binary thresholding in RePO also induces implicit data filtering, which we analyze in Section 3.3.
3.2 Relationship between SimPO and RePO
@@
Experimental setup. We adopt the experimental setup used in SimPO (Meng et al., 2024), with Llama3-8B and Gemma2-9B as our base models. For comparison, we use pre-trained instruction-tuned models as reference policies. To ensure consistency, we employ the same training datasets as in SimPO: princeton-nlp/llama3-ultrafeedback-armorm [^1] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm for Llama3-8B and princeton-nlp/gemma2-ultrafeedback-armorm [^2] https://huggingface.co/datasets/princeton-nlp/gemma2-ultrafeedback-armorm for Gemma2-9B. In all experiments involving SimPO, we set $\beta=10.0$ and $\gamma=0.4$ for Gemma2-9B and $\beta=10.0$ and $\gamma=0.3$ for Llama3-8B, unless stated otherwise.
@@
Evaluation benchmarks. We evaluate our models on two widely used open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024). These benchmarks assess the models’ conversational abilities across various queries. For AlpacaEval 2, we report both the length-controlled win rate (LC) and raw win rate (WR). For Arena-Hard, we report the win rate (WR).
Observation 1: Large $\beta$ enhances SimPO’s performance. We first experiment with SimPO by varying the parameter $\beta$ over the set {1.0, 3.0, 5.0, 10.0}, while keeping $\gamma$ fixed. As shown in Figure 3, increasing $\beta$ consistently improves SimPO’s performance,with convergence observed as $\beta$ grows larger
Observation 2: RePO achieves competitive performance. RePO achieves performance comparable to SimPO with a large $\beta$. As shown in Figure 3, RePO achieves win rates of 51.1% on Llama3-8B and 66.6% on Gemma2-9B, comparable to SimPO’s performance. This aligns with Lemma 3.1, which establishes that RePO can be interpreted as a limiting case of SimPO as $\beta\to\infty$.
Figure 4: Line plot illustrating the variation in RePO’s performance with different values of $\gamma$, measured by the AlpacaEval 2 LC Win Rate. The accompanying bar chart shows the mean implicit reward margin, $m_{\mathcal{D}}$, computed across all training pairs.
Observation 3: $\gamma$ is a crucial hyperparameter. In RePO, the gradients become zero when the implicit reward margin exceeds $\gamma$, indicating that the model $\pi_{\theta}$ is confident in its predictions for these data pairs. These pairs, considered well-separated, are excluded from updates, mitigating overfitting. Thus, $\gamma$ serves as a critical “cutoff point”. As shown in Figure 4, the performance of RePO is highly sensitive to $\gamma$. Increasing $\gamma$ initially improves performance, but beyond a certain threshold, performance degrades. Intuitively, A small $\gamma$ retains many zero-gradient data points, leading to under-filtering. Conversely, a large $\gamma$ causes many data points to have unit gradient weight, potentially leading to overfitting. Further discussion on $\gamma$’s role in data filtering is provided in subsequent sections.
@@
3.3 Data Filtering by RePO
Building upon the preceding discussion, we observe that adjusting the parameter $\gamma$ defines a “cutoff point” that determines which sample gradients are set to zero, effectively implementing data filtering. In this section, we analyze the training process of RePO through the lens of data filtering.
Observation 4: $\gamma$ controls the distribution of implicit reward margins. We also report $m_{\mathcal{D}}$ (cf. Equation 10), the mean of implicit reward margins across all training pairs. As shown in the bar chart of Figure 4, as $\gamma$ increases, $m_{\mathcal{D}}$ rises. However, alignment performance w.r.t. LC improves initially but declines once $\gamma$ exceeds a certain threshold. This suggests that while a larger $\gamma$ increases the separation between training pairs, excessive separation can introduce the risk of reward hacking (Amodei et al., 2016), leading to a performance drop.
Observation 5: Less data, better alignment. Figure 5 illustrates the distribution of the implicit reward margin $M_{\theta}$ across different training stages, alongside the ratio of zero gradients (i.e., when $M_{\theta}$ exceeds $\gamma$). As training progresses, the model’s ability to discriminate between winning and losing samples improves, resulting in a steady increase in both the implicit reward margin and the ratio of filtered data. Notably, the filtered data ratio rises from 48% to 58% between steps 200 and 400. Despite using only half of the samples for gradient updates, the model achieves optimal performance, indicating that RePO effectively utilizes a smaller, more informative subset of training data.

### v2

# RePO: Understanding Preference Learning Through ReLU-Based Optimization

@@ Abstract
Preference learning has become a common approach in various recent methods for aligning large language models with human values. These methods optimize the preference margin between chosen and rejected responses, subject to certain constraints for avoiding over-optimization. In this paper, we report surprising empirical findings that simple ReLU activation can learn meaningful alignments even using none of the following: (i) sigmoid-based gradient constraints, (ii) explicit regularization terms. Our experiments show that over-optimization does exist, but a threshold parameter $\gamma$ plays an essential role in preventing it by dynamically filtering training examples. We further provide theoretical analysis demonstrating that ReLU-based Preference Optimization (RePO) corresponds to the convex envelope of the 0-1 loss, establishing its fundamental soundness. Our “RePO” method achieves competitive or superior results compared to established preference optimization approaches. We hope this simple baseline will motivate researchers to rethink the fundamental mechanisms behind preference optimization for language model alignment.
@@ 1 Introduction
Recent years have witnessed significant advances in aligning large language models (LLMs) with human preferences [1, 2, 3, 4]. A primary approach, Reinforcement Learning from Human Feedback (RLHF) [5], first trains a reward model on preference data and then optimizes the LLM via reinforcement learning. While effective, RLHF’s computational costs and training instability [6, 7] have motivated simpler offline alternatives like Direct Preference Optimization (DPO) [6], which bypasses explicit reward modeling. Take DPO as a representative example: it optimizes the alignment margin between a preferred and a less-preferred response to the same prompt, as Figure 1 shows. The alignment of each response is quantified via an implicit reward, defined as the log-ratio of the predicted likelihoods under the policy model (i.e., the LLM being optimized) and a reference model (e.g., a fixed supervised fine-tuned (SFT) model).
A fundamental challenge in preference learning is over-optimization — where models excessively amplify reward margins between preferred and non-preferred responses, potentially degrading generation quality [8, 9, 10]. Several approaches have been developed to mitigate this issue. DPO [6] and SimPO [11] employ sigmoid weighting through log-sigmoid activation that diminishes gradients as reward margins increase, naturally preventing over-optimization. The $\beta$ parameter controls gradient distribution sharpness — larger values produce more binary-like gradients, as illustrated in Figure 2. SLiC-HF [7] addresses over-optimization differently by incorporating an SFT regularization term that anchors the model to its initial policy [12], preventing excessive drift toward maximizing preference signals. These mechanisms effectively balance preference optimization with generation quality preservation, forming the foundation of current preference learning approaches.
Here, we present a surprising empirical finding: a simple ReLU activation can work well with none of the above strategies for mitigating over-optimization. Our analysis reveals that as parameter $\beta$ in SimPO approaches infinity, its sigmoid weighting naturally converges to a binary thresholding mechanism — motivating our exploration of ReLU-based Preference Optimization (RePO). This mechanism uses a single ReLU function with only one hyperparameter $\gamma$, creating a clear decision boundary that selectively updates sample pairs with insufficient reward margins ($M_{\theta}<\gamma$) while filtering out well-separated pairs ($M_{\theta}\geq\gamma$). We illustrate this “RePO” method in Figure 1.
Figure 1: Comparing preference learning mechanisms. RePO employs a simpler binary thresholding mechanism than SimPO and DPO, as highlighted in the shaded box. Despite its simplicity, this mechanism achieves competitive results by naturally preventing over-optimization.
Figure 2: Gradient weighting functions of SimPO ($s_{\theta}$) and RePO ($\mathbb{I}(M_{\theta}<\gamma)$). As $\beta\to\infty$, $s_{\theta}$ converges to the binary indicator (red line), establishing RePO as the limit case of SimPO.
Thanks to its conceptual simplicity, RePO can serve as a hub that relates several existing methods. In essence, our method can be viewed as “SimPO without log-sigmoid” or “SLiC-HF without SFT regularization term”. Interestingly, RePO is related to each method by removing one of its core components. Even so, RePO effectively prevents over-optimization while performing competitively or better (cf. Figure 1).
We empirically show that over-optimization do exist, but ReLU activation with threshold $\gamma$ is critical to prevent such solutions. This implies that in over-optimization regimes, selecting which examples to learn from is more critical than determining how much to learn from each. The $\gamma$ threshold induces an emergent data filtering behavior, focusing dynamically on challenging samples relative to the model’s current capability. Our theoretical analysis reveals that RePO’s ReLU loss corresponds precisely to the convex envelope of the 0-1 loss (Theorem 4.2), explaining why such a simple mechanism is so effective.
Our simple baseline suggests that the ReLU activation with a proper threshold $\gamma$ can be an essential reason for the common success of related methods. We believe this work’s significance lies in revealing how preference learning principles may be simpler than previously thought. By questioning conventional wisdom about necessary components, we hope to motivate researchers to reconsider the fundamental mechanisms behind preference optimization.
@@ 2 Preliminaries
Directed Preference Optimization (DPO). DPO [6] stands out as a leading method for offline preference optimization by eliminating the need for an explicit reward model. Instead, it reformulates the reward $r(x,y)$ as a closed-form expression based on policy ratios:
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right)\right],$ (2)
@@
Sequence Likelihood Calibration (SLiC-HF). SLiC-HF [7] advances preference optimization with two key innovations: (1) it employs a sequence-level calibration loss that contrasts the log-probability difference between preferred and dispreferred responses using a margin $\gamma$, and (2) it integrates a regularization term to prevent divergence from the SFT policy, avoiding the need for an explicit KL penalty. The SLiC-HF loss function is defined as:
$\mathcal{L}_{\text{SLiC-HF}}(\pi_{\theta})=$ $\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\Big[\text{ReLU}\Big(-\Big(\log{\pi_{\theta}(y_{w}\mid x)}-\log\pi_{\theta}(y_{l}\mid x)-\gamma\Big)\Big)-\lambda\log{\pi_{\theta}(y_{w}\mid x)}\Big].$ (3)
Simple Preference Optimization (SimPO). SimPO [11] advances preference optimization with two key innovations: (1) it normalizes the reward by the length of the response, calculating the average log-probability per token for a response under the policy $\pi_{\theta}$, and (2) it incorporates a target reward margin $\gamma$ to ensure that the reward difference between the preferred and less preferred responses exceeds this margin. The SimPO loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|}-\gamma\right)\right)\right],$ (4)
@@
3 Exploring Simple ReLU Activation in Preference Learning
In this section, we explore what makes a simple ReLU activation function effective for preference learning. We first examine the surprising relationship between ReLU activation and sigmoid weighting through empirical experiments. Then, we investigate the key properties that emerge from this simple mechanism, specifically through the lens of gradient behavior, data filtering patterns, and over-optimization control.
3.1 Examining ReLU-based Preference Optimization
Simplification exploration. Our exploration began by questioning whether log-sigmoid activation or SFT regularization are truly necessary for mitigating over-optimization. We simplified the SimPO loss function through two key modifications: (i) removing the hyperparameter $\beta$, and (ii) replacing the log-sigmoid function with a ReLU activation.
We adopt the length-normalized implicit reward margin $M_{\theta}$ (as introduced in SimPO [11]):
@@ 3.1 Examining ReLU-based Preference Optimization
which quantifies the policy’s preference between responses. Using $M_{\theta}$, we examine a loss function with the following form:
@@
Gradient behavior investigation. We examine the gradient dynamics of RePO and SimPO to reveal how our simplified approach addresses over-optimization:
@@
where $s_{\theta}=\sigma(\beta(-M_{\theta}+\gamma))$ is SimPO’s sigmoid weighting function. The terms $\nabla_{\theta,y_{w}}=\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)$ and $\nabla_{\theta,y_{l}}=\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)$ correspond to the gradients that increase the probability of the “winning” response $y_{w}$ and decrease the probability of the “losing” response $y_{l}$, respectively. The scaling factor $\beta$ in Equation 7 linearly amplifies gradient magnitudes but does not alter the relative update directions in adaptive optimizers like Adam [13], as the momentum terms automatically normalize scale variations. We therefore omit $\beta$ in Figure 2 for clearer visualization of the weighting function shapes.
The key insight is that RePO’s ReLU-based gradient (Equation 8) applies uniform updates only to pairs with $M_{\theta}<\gamma$, while SimPO’s gradient (Equation 7) uses continuous $\beta$-scaled weights. Figure 2 visualizes this difference, showing RePO as the limiting case of SimPO as $\beta\to\infty$.
@@
Please check Appendix for all proofs. Lemma 3.1 establishes RePO as the asymptotic limit of SimPO with large $\beta$, explaining two key advantages we will demonstrate in Section 3.2: comparable performance without $\beta$ tuning complexity, and an effective binary thresholding mechanism that induces implicit data filtering for controlling over-optimization.
3.2 Empirical Study
@@
Experimental setup. We evaluate this approach using SimPO’s experimental setup [11] with Llama3-8B and Gemma2-9B models (Instruct setup). For consistency, we use the same training datasets as SimPO: princeton-nlp/llama3-ultrafeedback-armorm for Llama3-8B and princeton-nlp/gemma2-ultrafeedback-armorm for Gemma2-9B. For all SimPO experiments, we set $\beta=10.0$ and $\gamma=0.4$ for Gemma2-9B and $\beta=10.0$ and $\gamma=0.3$ for Llama3-8B, unless otherwise specified. We track optimization progress using two reward margin metrics:
$m_{\text{batch}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{B}}[M_{\theta}],\quad m_{\mathcal{D}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[M_{\theta}],$ (10)
measuring response separation within each batch ($m_{\text{batch}}$) and across the entire training set ($m_{\mathcal{D}}$).
@@
Evaluation benchmarks. We evaluate on two established benchmarks for open-ended generation: AlpacaEval 2 [14] (measuring instruction-following quality against GPT-4) and Arena-Hard [15] (testing complex reasoning). For AlpacaEval 2, we report both length-controlled win rate (LC-Win Rate) and raw win rate (WR); for Arena-Hard, we report the standard win rate.
Observation 1: Large $\beta$ enhances SimPO’s performance when paired with appropriate $\gamma$. We systematically evaluate SimPO across varying values of $\beta\in\{1.0,3.0,5.0,10.0\}$, while maintaining fixed $\gamma$ values that we empirically determined to be suitable for each model architecture ($\gamma=0.4$ for Gemma2-9B and $\gamma=0.3$ for Llama3-8B). As shown in Figure 3, increasing $\beta$ leads to consistent performance improvements across all evaluation metrics, with diminishing returns observed beyond $\beta=5.0$. These findings align with observations in the SimPO paper [^1] In their official repository, the authors note: “SimPO requires a much larger $\beta$ than DPO… In many cases, an even larger (e.g., 10) could yield better results.”.
@@
Figure 5: Line plot of RePO performance (AlpacaEval 2 LC Win Rate) and bar chart of mean reward margins ($m_{\mathcal{D}}$) across varying $\gamma$ values. See Appendix D.3 for details.
Observation 2: RePO matches high-$\beta$ SimPO. RePO achieves performance comparable to SimPO with a large $\beta$. As shown in Figure 3, RePO achieves win rates of 51.1% on Llama3-8B and 66.6% on Gemma2-9B, comparable to SimPO’s performance. This aligns with Lemma 3.1, which establishes that RePO can be interpreted as a limiting case of SimPO as $\beta\to\infty$.
Observation 3: $\gamma$ threshold creates a natural alignment-optimization tradeoff. Our experiments track the mean implicit reward margin $m_{\mathcal{D}}$ (cf. Equation 10) across training pairs. As Figure 5 illustrates, increasing $\gamma$ directly elevates $m_{\mathcal{D}}$ while performance follows an inverted U-shaped pattern — improving initially but declining beyond a critical threshold. In RePO, gradients vanish when the implicit reward margin exceeds $\gamma$, effectively filtering out well-separated pairs from updates. This mechanism creates a fundamental tradeoff: small $\gamma$ values retain excessive zero-gradient samples causing under-filtering, while large $\gamma$ values force updates on most samples, potentially leading to over-optimization [8] and ultimately degrading performance.
Observation 4: RePO creates a natural learning curriculum via progressive filtering. Figure 4 reveals an unexpected pattern in how the distribution of implicit reward margins $M_{\theta}$ evolves throughout training. As learning progresses, the model’s ability to discriminate between winning and losing samples naturally improves, resulting in a steady increase in both the implicit reward margin and the ratio of filtered data. Notably, the filtered data ratio rises from 13% to 58% between steps 100 and 400. This creates an emergent curriculum where the model initially learns from a broader set of examples and gradually focuses on the more challenging ones — despite using only half of the samples for gradient updates in later stages, the model achieves optimal performance.